# LLM-Deliberation: Evaluating LLMs with Interactive Multi-Agent Negotiation Games

### v1

@@ 4.2 Prompting Strategy: Incentives and Structure
Incentives: One out. In this game variant, one party is instructed to sabotage the negotiation. Thus, it is told to “not care about being fair or accommodating others”. The prompts instruct the agent to either sabotage the negotiation or at least maximize its own score as much as possible if the negotiation seems likely to succeed. This player gets a higher score if no deal is achieved. To provide a mechanism of how sabotaging can be possible, we instruct the agent to “isolate one party by pushing for deals that you think they will oppose, but others might support”.
Structure. We use structured CoT to instruct agents to plan their answers and show intermediate calculations in a “Scratchpad”. We instruct the agent to indicate the scratchpad by special tokens, such that we remove them from the shared negotiation history that is public to all agents. In the scratchpad, the agent collects observations and information (e.g., prompting it to “think about what others may prefer”, “observe others’ interactions”), then explores possible next moves that satisfy its goals (denoted by the incentive, e.g., cooperation or greediness, etc.). These two steps are inspired by Gandhi et al. (2023b) and aim to explicitly decompose the task into smaller ones. To limit the context window size, the agent does not see its scratchpad from previous rounds. We observed that this might cause a (metaphorical) “discontinuation of thoughts”, e.g., an agent’s utterance might contain references to actions it can explore the next time (e.g., “I will propose $a_{1}$ first, and if others disagree, I can compromise to $a_{2}$”), or the agent may consistently propose the same deal each round. Longer-term planning is also integral to how humans negotiate (LSB, ). To simulate this, the agent is prompted to generate a plan after its final answer of actions it can take the next time. At the next agent’s turn, the agent is fed its previous “plan” if it exists. Agents do not see others’ plans. We show more ablations on the structure in our experiments.
Figure 3: Example from a negotiation session. The agent takes the previous interactions appended to its initial prompts. The prompt incentivized the agent to cooperate and is structured as observation, exploration, and planning steps.
@@ 5 Experiments and Evaluation
In this section, we present our experimental evaluation. We first describe the experimental setup and demonstrate an ablation of the prompt structure and models’ comparison. We use the base game and the “All in - cooperative” variant as a baseline. Next, we show the performance on different games and discuss how the game’s difficulty can be tuned. Finally, we illustrate other game variants where one agent has a different incentive while the others are cooperative.
5.1 Experimental Setup
@@
We used 24 rounds, with 4 consecutive random ordering of the 6 agents (i.e., each agent is prompted 4 times). We use a history window of the last 6 interactions. We compared GPT-4 vs. GPT-3.5 due to their high performance on LLMs-as-Agent benchmarks (Liu et al., 2023) (the gpt-4-0613 and gpt-3.5-turbo-16k-0613 models’ snapshots). For reproducibility, we used temperature 0 for sampling in all our experiments. Models are instructed to enclose the scratchpad, the final answer, the plan, and any suggested deals between specific tags to enable automatic parsing and calculation of deals’ scores. We ran each experiment 20 times (with a different random order of agents’ interactions) to compute the success rate in reaching an agreement and the average of scores over the multiple runs. Specifically, we propose the following metrics to quantify the performance:
  • [leftmargin=*]
  • Final success: Rate of games with a successful deal made by $p_{1}$ at the end of negotiation, $\pi_{p_{1}}^{(r+1)}\in\Pi_{\text{pass}}$. Using $p_{1}$ as an anchor of success is the most challenging because its preferences are at odds with most parties. We also compute the rate of final deals that lead to a 6-way agreement (i.e., all).
@@
  • Any success: Rate of games with a successful deal made by $p_{1}$ at any round, $\pi_{p_{1}}^{(t)}\in\Pi_{\text{pass}}$.
  • Own score: For an agent $p_{i}$, we calculate the score of the deal it proposes w.r.t. itself, $S_{p_{i}}(\pi_{p_{i}}^{(t)})$. This is a “local view” of the agent’s actions (i.e., its proposed deals). This is relevant to measuring whether agents are proposing deals that are less ideal for them (i.e., compromising).
  • Collective score: For an agent $p_{i}$, we calculate the average score of all agents given its deals $\pi_{p_{i}}^{(t)}$,
    $\frac{1}{|P|}\sum\limits_{p_{j}\in P}S_{p_{j}}(\pi_{p_{i}}^{(t)}),$
    this is an “oracle view” of the agent’s actions w.r.t. all agents; the agent cannot observe the value of the collective score. This can measure whether agents are approaching a common agreement.
@@
  • Wrong deals: Rate of deals $\pi_{p_{i}}^{(t)}$ with own score $S_{p_{i}}(\pi_{p_{i}}^{(t)})$ that is less than the corresponding minimum threshold of $p_{i}$. This measures whether models are following the game’s rules and performing correct calculations of deals.
5.2 Probing Models via an Ablation of Prompts’ Structure
We study different variants of the prompt structure provided to the model at each round, while keeping the initial prompts fixed. We first show the “no-CoT” performance. Next, we evaluate the role of the CoT “planning” stage and different alternatives of the CoT “observation” and “exploration” stages that break them down into steps to provide more or less guidance.
CoT: Observation and Information Collection: This involves a “previous deals’ calculation” step in which we prompt agents to calculate their scores of each deal that was proposed in the current history window. Then, we follow this with “inferring others’ preferences”, instructing agents to “think about others’ preferences”. Our ablation removes the first or both steps.
CoT: Exploration: We prompt agents to perform candidate generation by finding 3 different deals that are higher than their minimum thresholds while considering” other preferences and their previous plans, if any. Then, as a last step, we prompt agents to make a final proposal selection that is likely to achieve their goal (depending on the game variant: balancing between their preferences and others, maximizing their scores, or sabotaging the negotiation).
As shown in Table 1, GPT-4 consistently achieved significantly higher performance. We summarize our main findings as follows:
@@ 5.2 Probing Models via an Ablation of Prompts’ Structure
Model CoT: Observation CoT: Exploration CoT: Planning Final normal-$\uparrow$ Any normal-$\uparrow$ Wrong normal-$\downarrow$
@@
Table 1: Prompt structure ablation study, shown in rows. Yellow markers indicate changes in the experiment compared to the previous row. We break the observation phase into previous deals’ calculation and inferring others’ preferences steps, and the exploration phase into candidate generation and final deal selection steps. The metrics (%) show the rate of games with a successful deal at the end of the negotiation (Final) with at least a 5-way agreement or a 6-way agreement, the rate of games with a successful deal at any round during the negotiation (Any), and the rate of deals suggested by agents that violate their own minimum threshold (Wrong).
(a) Best.
(b) “No planning”.
(c) “No others”.
(d) Worst (all steps).
Figure 4: The “own score” and “collective score” of $p_{1}$’s deals over rounds of GPT-4 experiments in Table 1. In (a), the best prompt, the “own score” continues to decrease (but rarely below the minimum threshold), and the “collective score” continues to increase. In (b) and (c), the scores saturate. In (d), the prompt with all steps, the “own score” is relatively higher; agents consistently explored and proposed deals that are more ideal to them rather than adapting to observations.
1) Evaluation: calculations. GPT-3.5 agents often propose deals that are less than their minimum thresholds (the “wrong deals”). This percentage is almost negligible in GPT-4 agents, especially when using CoT. When inspecting the CoT, we observed that GPT-4 agents mostly correctly calculate and compare the value of these wrong deals to their thresholds, but their CoTs sometimes contain phrases such as, “this deal is slightly less than my minimum threshold, but I am willing to compromise if this would lead to an agreement”, which, however, may indicate a lack of exploration.
2) Evaluation: CoT variations. Even when using no CoT, GPT-4’s performance is relatively high. Interestingly, guiding the reasoning with more intermediate steps seems to help with GPT-3.5 agents but hurt the performance of GPT-4. We observed that when explicitly prompted to explore feasible solutions, GPT-4 agents were biased towards generating and selecting deals that scored consistently higher for them, rather than adaptively finding deals that integrate the current observations or that are conditioned on cooperation. We also found that the planning step is important, especially to reach a deal at the end. A similar observation can be made regarding the others’ preferences inference step. To demonstrate these conclusions, we show in Figure 4 the progression of $p_{1}$’s deals over rounds.
@@
3) Evaluation: ToM and qualitative examples. Additionally, to test whether models can explicitly infer the preferences of others, we prompted each agent to provide a “best guess” of each party’s preferred sub-option under each issue. Each agent sees its own initial instructions only. This is done before any interaction (to test commonsense reasoning based on the game’s semantics without observations from other agents). GPT-4 models scored 61% in correctly matching the ground truth preferences of sub-options, vs. 42% by GPT-3.5 (averaged over all agents and multiple trials). GPT-4 models frequently correctly assigned neutral values for issues with no clear associations (e.g., “the Green Alliance might not have any preference on employment distribution”), and made a distinction between $P_{\text{oppose}}$ and $P_{\text{benefit}}$ regarding implicit preference entailment (e.g., “they might want to limit/ensure the project’s success by requesting less/more funding”) even though this distinction was not provided in the initial prompt. We show an example of $p_{1}$’s CoT in Figure 3. In this example, the agent (GPT-4) iteratively alters its suggestion to accommodate $p_{2}$ (after a correct inference of its preference) and to meet its own score. In contrast, GPT-3.5 agents (see Appendix D) may leak their secret scores in their public answer and argue for deals because they have high scores (indicating a lack of ToM-related reasoning). Given the superior performance of GPT-4 agents, we focus the rest of the experiments on this model to further evaluate its behavior.
Game Final normal-$\uparrow$ Any normal-$\uparrow$ Wrong normal-$\downarrow$
@@
Base (55/12) 81 33 100 1.4
@@
Base (30/4) 65 25  85 0.7
Base (17/2) 30  5  70 2.0
Table 2: Performance (%) on new games and difficult levels of the base game. Numbers between brackets denote the feasible 5-way and 6-way agreements, respectively.
5.3 Performance on Other Games
Next, we evaluate GPT-4 (with the best prompting strategy) on other games, as presented in Table 2. We rewrite the base game by prompting GPT-4 to change the entity names while maintaining semantic relationships. We also swapped issue names (e.g., $A\rightarrow D$, etc.). This is intended to test the robustness against semantically similar changes and to rule out memorization. As shown, the performance on the base and rewritten games is comparable. Furthermore, agents continue to perform relatively well on the newly created games, especially considering the “Any success” metric. They, however, have varying levels of performance when considering the final agreement. While all games have a comparable number of feasible solutions, games 1 and 2 can be relatively harder because they have non-spare scores (i.e., all agents have preferences on almost all issues). This might require more fine granularity when proposing deals; from the perspective of one agent, deals with comparable or even the same scores might have a highly fluctuating number of agreeing parties. Therefore, to match the base game, we designed the third game to have more sparse scores, which indeed scored similarly regarding the final deal metric (the 6-way agreement is higher because the set of 6-way feasible solutions is larger). More analysis of the games’ difficulty is in Appendix A, where we also show the progression of $p_{1}$’s deals over rounds, which follows a similar trend to 3(a).
@@ 5.4 Tuning the Game Difficulty
Besides designing differently sparse games, the difficulty of games may also be easily tuned by changing the minimum thresholds of agents and re-running the simulations while keeping the issues, priorities, and scores fixed. This can create new alterations to games and help adapt our evaluations to future advanced models. This is important since we witness a saturation of older benchmarks with the release of more powerful models, which may, in turn, misleadingly imply that a particular task has been essentially solved. However, models still fail with sometimes trivial alterations (Ullman, 2023). These difficult variations can also help foster future research; performance drops when we decrease the set of feasible solutions (the last part in Table 2), indicating that improved paradigms in communication and exploration (Hao et al., 2023) may be needed.
5.5 Other Game Variants: Greedy and Saboteur Agents
So far, all agents are explicitly incentivized to compromise toward a common agreement. However, when agents interact in the real world with other agents or humans, they might face adversarial or greedy players. To simulate such setups, we vary the incentive of one player while keeping the others fixed, creating the two previously mentioned variants, “All in - one greedy” and “One out”. For “One out”, we further have an untargeted variant, where the agent is not given specific instructions on which agent to “attack” (i.e., propose deals against), and a targeted variant that specifies the target agent. We aim to answer two main questions:
(a) All in.
(b) All in - one greedy.
(c) One out (untargeted).
(d) One out (targeted).
Figure 5: The “own score” and “collective score” of the same agent $p_{i}\in P_{\text{const}}$ in the cooperative, greedy, and sabotaging (untargeted and targeted) games. $p_{i}$ is suggesting the deals at its turn. We assign another agent $p_{v}$ as the target. $p_{i}$’s actions are consistent with its assigned incentives (compromising, maximizing score, pushing for non-agreeable deals, and targeting an agent, respectively).
1) Are agents’ actions consistent with their assigned incentives? Agents’ specific actions can be modulated by high-level incentives that do not specify particular deals. We calculate the “own score” and “collective score” of the same agent assigned with the cooperative, greedy, and sabotaging incentives, shown in Figure 5. In the cooperative game, the “own score” is the lowest, while the “collective score” is high, with a tight gap. In the greedy game, the “own score” is higher, but the agent is still finding deals that might be agreeable (i.e., indicated by a relatively high “collective score”). In the sabotaging game, the “own score” is also high, but the agent’s suggested deals give a low “collective score”. Moreover, in the targeted version, the target’s score is lower compared to the untargeted case. It is important to note here that the agent cannot see others’ scores.
2) What are the effects on other players and the success of the negotiation? We show the final success rate in Table 3. Both variants decrease the performance compared to the cooperative game; the greedy/saboteur agents’ actions affected the group.
Variant Final normal-$\uparrow$

### v2

@@ 3.2 Compromising, Greedy, and Adversarial Games
Adversarial game. Here, one party is instructed to sabotage the negotiation or at least maximize its own score as much as possible if the negotiation seems likely to succeed. This player gets a higher score if no deal is achieved. This is, their BATNA is higher than 100 (the maximum achievable score). To provide a mechanism for sabotaging, we instruct the agent to “isolate one party by pushing for deals that you think they will oppose, but others might support”. We conduct two experiments: one where we specify the victim/target agent $p_{v}$ (targeted) and one where the agent autonomously picks one (untargeted). Similar to the greedy game, $w_{p_{i}}\gg\text{max}(\{w_{p_{j}}|\;p_{j}\in P\backslash\{p_{i}\}\})$. In addition, $w_{p_{v}}<0$ (to minimize the target’s score). This would result in a lower average score for the group.
Natural language incentives. We verbalize these variants as high-level “incentives” given to the model in the initial and round prompts; e.g., compromising agents are instructed to aim for a balanced deal, accommodate other parties, etc. Adversarial agents are instructed to “not care about being fair or accommodating others”, etc. However, we do not instruct agents on which deals to propose.
Assumptions. In all variants, agents are not prompted with any information about other players’ incentives. In the adversarial variant, a successful deal has to satisfy the thresholds of the other 5 parties. We introduce only one adversary to have a similar success condition across variants.
3.3 A Baseline Prompting Solution Framework
We use structured Chain-of-Thought [44] to enable agents to decompose the task, plan their answers, and show intermediate calculations in a secret “scratchpad”. We use the following structure:
CoT: Observation. The agent first should collect observations and information from the ongoing history. This involves a “previous deals’ calculation” step in which we prompt agents to calculate their scores for each deal that was proposed in the current history window. Then, we follow this with an instruction to “infer others’ preferences”. We remove one or both steps in our ablation.
CoT: Exploration. Next, agents should explore possible moves by “generating candidates”, i.e., 3 potential deals that are higher than their thresholds, then “selecting a final deal” that is likely to achieve their respective goal. Our ablation removes the first step.
CoT: Planning. Planning is integral to how humans negotiate [18]. We observed agents’ utterances may contain references to actions they can explore the next time (e.g., “I will propose $a_{1}$ first, then, I can compromise to $a_{2}$”). Without long-term planning and a limited shared history, the agent might propose similar deals each round. Therefore, as long as the agent has a next turn, we instruct it to generate a secret plan of possible next actions. At the next turn, the agent is fed its respective previous “plan” appended to the round’s prompt $C_{p_{i}}^{(t)}$. Agents’ output in eqn. 1 can thus be broken down as:
$O_{p_{i}}^{(t)}\coloneqq\left\{\begin{array}[]{ll}\left[\sigma_{p_{i}}^{(t)},\alpha_{p_{i}}^{(t)},\rho_{p_{i}}^{(t)}\right]&\text{if next(}p_{i}\text{) = }\texttt{True}\\ \left[\sigma_{p_{i}}^{(t)},\alpha_{p_{i}}^{(t)}\right]&\text{otherwise,}\\ \end{array}\right.$ (5)
, $\sigma_{p_{i}}^{(t)}$ is the scratchpad, $\alpha_{p_{i}}^{(t)}$ is the public answer, and $\rho_{p_{i}}^{(t)}$ is the plan.
@@ 4 Experiments and Evaluation
We first describe our setup and show the ablation study and models’ comparison. Next, we show the performance of other games and the greedy and adversarial variants.
4.1 Experimental Setup and Evaluation Metrics
@@
We used 24 rounds, with 4 consecutive random ordering of the 6 agents and a history window of the last 6 interactions. We test on GPT-4, GPT-3.5, Gemini Pro, Llama2 13b and 70b Chat, Llama3 70b Chat, and Mixtral 8x7B. For reproducibility, we used a sampling temperature of 0. Models are instructed to indicate deals, scratchpads, public answers, and plans by specific tags to enable automatic parsing and calculation of deals’ scores. We ran each experiment 20 times (with a random order of agents) to compute the average performance. Specifically, we propose the following metrics:
@@
Final success. Rate of games with a successful final deal (made by $p_{1}$ at the end of the negotiation), i.e., $\pi_{p_{1}}^{(R+1)}\in\Pi_{\text{pass}}$. We measure both 5-way and 6-way agreement rates.
Any success. Rate of games with a successful deal by $p_{1}$ at any time; $\pi_{p_{1}}^{(t)}\in\Pi_{\text{pass}}$ is True for any $t$.
Own score. We calculate $p_{i}$’s scores of its proposed deals w.r.t. itself: $S_{p_{i}}(\pi_{p_{i}}^{(t)})$. This is a “local view” of the agent’s actions and helps measure if/how agents are aligned with their roles.
Collective score. For an agent $p_{i}$, we calculate the average score of all agents given its deals: $\frac{1}{|P|}\sum\limits_{p_{j}\in P}S_{p_{j}}(\pi_{p_{i}}^{(t)})$. This is an “oracle view” of the agent’s actions w.r.t. others, which $p_{i}$ cannot observe. This measures whether agents make correct inferences about others’ goals and take actions that are likely to achieve their goals (e.g., agreement, sabotaging).
@@
Wrong deals. Rate of deals with “own score” less than the corresponding minimum threshold of the agent: $S_{p_{i}}(\pi_{p_{i}}^{(t)})<\tau_{p_{i}}$. This measures whether models are performing correct calculations of deals.
Score leakage ratio. Agents were instructed not to reveal information about scores. This is usually a critically needed behavior in practical negotiation setups. This also broadly measures the trustworthiness of models in following instructions and keeping in-context confidential information [7], a task that is also related to ToM [24]. We use GPT-4 as a judge to verify whether public answers contain any mention of scores or thresholds, and we compute the ratio of answers with leaked scores.
@@
Model row no. CoT: Observation CoT: Exploration CoT: Planning Final (%) $\uparrow$ Any (%) $\uparrow$ Wrong (%) $\downarrow$
@@
Table 1: Prompt structure ablation study. Yellow markers indicate changes in the experiment compared to the previous row. The prompt structure is: score calculation of previous deals in the public history, inferring others’ preferences, candidate generation, final deal selection, and planning.
(a) Best (row 5).
(b) “No plan” (row 4).
(c) “No others” (row 6).
(d) All steps (row 2).
Figure 3: $p_{1}$’s deals progression over rounds of GPT-4 experiments in Table 1. In (a), the “own score” decreases, and the “collective score” increases, indicating more agreement. In (b) and (c), they stop improving and saturate during the final rounds. In (d), agents proposed deals that are more ideal to them and which do not increase the collective score, lowering the success in reaching an agreement.
4.2 Ablation of Prompts’ Structure
As discussed in Section 3.3, we study variants of the prompt structure given to agents at each round $C_{p_{i}}^{(t)}$. We remove the planning stage and vary the CoT “observation” and “exploration” stages. We also evaluate the no-CoT performance. We perform an ablation study on GPT-3.5 and GPT-4 and later test on the other models with the best-found configuration. Rows in Table 1 show these experiments, averaged over runs. Figure 3 shows the progression of $p_{1}$’s deals over rounds to visualize whether (and by how much) $p_{1}$ is successfully reaching agreement in the GPT-4 experiments. Our analysis, depicted next, aims to reveal which skills are needed to reach success.
Arithmetic calculations. GPT-3.5 agents often propose deals that are less than their minimum thresholds (indicated by a higher value of the “wrong deals” metric). This is almost negligible in GPT-4 agents, especially when using CoT. In addition to computing the “wrong deals”, tracking agents’ deals can also evaluate how well agents follow instructions and are aligned with their assigned payoffs and negotiation roles, rather than following pretraining biases that would make some options more ideal; we show in Appendix B a histogram in which GPT-4 agents advocate or oppose strong environmental protection measures in consistency with their respective payoffs.
Figure 4: Example from GPT-4 simulation. The agent takes the interaction history along with its initial prompt and instructions that incentivize it to cooperate, which are structured as observation, exploration, and planning steps. The agent here autonomously and iteratively adjusts its suggestions.
@@ 4.2 Ablation of Prompts’ Structure
ToM. In Table 1, we show that instructing models to infer others’ preferences increases the success rate (indicated by the drop in rows 6 and 10). To test if models can explicitly infer the preferences of others, we further prompted each agent to provide a “best guess” of each party’s preferred sub-option under each issue. Each agent sees only its own initial instructions $C_{p_{i}}^{0}$ before interaction (to test commonsense reasoning based on the game’s semantics, without observations from other agents). GPT-4 models scored 61% in correctly matching the ground truth preferences of sub-options, vs. 42% by GPT-3.5 (averaged over all agents). GPT-4 models frequently correctly assigned neutral values for issues with no clear associations (e.g., “the Green Alliance might not have any preference on employment distribution”) and made a distinction between $P_{\text{oppose}}$ and $P_{\text{benefit}}$ regarding implicit preference entailment (e.g., “they might want to limit/ensure the project’s success by requesting less/more funding”) even though this distinction was not provided in the initial prompt. In contrast, GPT-3.5 agents often leak their secret scores in their public answer and argue for deals because they have high scores, indicating a lack of ToM-related reasoning (see Appendix H and Table 3 next).
Adaptation and Exploration. GPT-3.5 agents benefited from instructions to explore feasible solutions (row 11), possibly due to improvements in calculations. However, when doing so with GPT-4, agents were biased towards generating deals and selecting the ones from the history that scored higher (see 3(d)). Without this step, GPT-4 agents were more likely to find deals that adapt to other parties (see row 2 vs. row 3). We show an example of $p_{1}$’s CoT in Figure 4 in which the GPT-4 agent iteratively alters its suggestion to accommodate $p_{2}$ (after a correct inference of its preference) and to meet its own threshold. However, we still observe a lack of exploration when the agent compensated by over-increasing its score in one issue instead of finding a balanced proposal.
Planning. This step was important to reach a final successful deal (row 4); without it, agents’ suggestions may saturate and no longer increase the collective score (3(b)).
4.3 Mixed Population
Models Final $\uparrow$
All GPT-4 81
All GPT-3.5 20
$p_{1}$ is GPT-3.5 50
$P_{\text{benefit}}$ are GPT-3.5 62
Table 2: Success (%) with a mixed population of models.
As future multi-agent systems might have asymmetrical individual units, we next study a mixed population of GPT-4 and GPT-3.5. Since the game involves cooperation, less capable models may result in lower success for the entire group. We show experiments in Table 2 with details in Appendix C. The main results are 1) including GPT-3.5 drops the success for the entire group, with the highest drop when $p_{1}$ is GPT-3.5; everyone is worse off, 2) GPT-3.5 agents can get lower scores than their counterparts in the ‘all GPT-4’ experiment.
Model Final $\uparrow$ Any $\uparrow$ Wrong $\downarrow$ Leaked $\downarrow$
@@ 4.3 Mixed Population
GPT-4 81 33 100 1.4 0
GPT-3.5 20 8 33 19 25
Llama2-13b 57 10 82 16 14
Llama2-70b 76 19 95 11 22
Llama3-70b 60 21 100 4 2
Gemini Pro 45 0 70 13 6
Mixtral 8x7B 65 17 95 11 12
Table 3: Performance (%) of other models.
4.4 Other Open-Source Models
We use the best prompt template from our ablation (on GPT-4) to test other models. We excluded Mistral 7b [12] and Llama2 7b as they did not follow the basic formatting of the game. As shown in Table 3, open-source models perform worse than GPT-4 but better than GPT-3.5. Llama3 70b comes close to GPT-4 considering agreement success, correct calculations, and not revealing scores. Other models are especially worse in calculation and keeping confidential scores (higher wrong deals and leaked scores ratios). I.e., our benchmark is already challenging for many SoTA models, and as shown next, its difficulty can be increased to test future models. Due to its higher performance, we perform the rest of our analysis on GPT-4.
4.5 Performance on Other Games
To test robustness against semantically similar changes, we rewrite the base game by prompting GPT-4 to change the entities and issue names while maintaining semantic relationships. As shown in Table 4, the performance on the base and rewritten games is comparable. Also, agents perform relatively well on the new games (created from scratch) with varying levels of success. While all games have a comparable number of feasible solutions, games 1 and 2 can be more competitive as they have non-sparse scores (i.e., all agents have preferences on almost all issues). This might require more fine granularity when proposing deals; from the perspective of one agent, deals with comparable or even the same scores might have a highly fluctuating number of agreeing parties. Therefore, to match the base game, we designed game 3 to have more sparse scores, which indeed scored similarly w.r.t. the final deal metric. More analysis of the games’ difficulty is in Appendix D. In summary, our benchmark has diverse and easily tunable difficulty levels to test future advanced models.
Game Final $\uparrow$ Any $\uparrow$
5/6-way 6-way
Base (55/12) 81 33 100
@@ 4.5 Performance on Other Games
Base (30/4) 65 25 85
Base (17/2) 30 5 70
Table 4: Success (%) of GPT-4 on new games and difficult levels of the base game. (#/#) are 5-way and 6-way deals, respectively.
@@ 4.6 Tuning the Game Difficulty
Besides designing diverse games, the difficulty of games can be easily tuned by changing agents’ minimum thresholds and re-running the simulation while keeping everything else fixed. This is critical since we witness a saturation of older benchmarks with the release of powerful models and training data contamination [42, 15]. Our evolving benchmark can help foster future research as there is still ample room for improvement; success drops when we decrease the set of feasible solutions (the last part in Table 4), indicating that advanced paradigms in communication, exploration, and planning can be incorporated. In addition, decreasing the number of players can be used to create easier games, as shown in our experiment in Appendix E, in which simulations with fewer agents have higher all-way agreement rates. This further motivates our multi-agent setup as it results in a more challenging environment.
(a) Compromising.
(b) Greedy.
(c) Adv. (untargeted).
(d) Adv. (targeted).
Figure 5: The “own score” and “collective score” of the same agent’s deals, $p_{i}\in P_{\text{const}}$, in the different variants. Another agent $p_{v}$ is the target in the targeted adversarial variant. $p_{i}$’s actions are consistent with its assigned incentives.
4.7 Greedy and Adversarial Variants
We now study the other variants discussed in Section 3.2 and aim to answer two main questions:
1) Are agents’ actions consistent with their high-level incentives? We calculate the “own score” and “collective score” of the same agent assigned with the different incentives, as shown in Figure 5. In the compromising variant, the “own score” is the lowest, while the “collective score” is high. In the greedy variant, the “own score” is higher, but the agent is still finding deals that might be agreeable (i.e., indicated by a relatively high “collective score”). In the adversarial variant, the “own score” is also high, but the agent’s suggested deals give a low “collective score”. In the targeted version, the target’s score is lower compared to the untargeted case. It is important to note that the agent cannot see others’ scores and that instructions never included what specific deals to propose. While GPT-4 mapped these incentives to plausible corresponding deals, GPT-3.5 failed to do so (see Figure 19), indicating that this is a non-trivial task.
Variant Final (%) $\uparrow$